OpenStudy (anonymous):

Assume f(x)=4x+8 and g(x)=5. What is the value of (f+g)(2)

OpenStudy (anonymous):

(f+g)(x)=4x+8+5=4x+13 , x=2 =>(f+g)(2)=4*2+13=21